# Profinite rigidity witnessed by Dehn fillings of cusped hyperbolic 3-manifolds

### v2

@@ 5.1. Detecting prime decomposition and asphericity
Let $M$ and $N$ be compact, orientable 3-manifolds with empty or toral boundary. The prime decomposition of $M$ and $N$ yields the following connected sum
@@
We Dehn fill the compressible boundary components of $M$ along their null-homotopic slopes, and obtain a boundary incompressible 3-manifold
@@
To be explicit, up to a choice of conjugacy representatives, $\pi_{1}M_{i}$ can be viewed as a free factor of ${\pi_{1}M^{+}}$. [41, Proposition 9.1.8] implies that $\pi_{1}M^{+}$ induces the full profinite topology on $\pi_{1}M_{i}$, so the homomorphism $\widehat{\pi_{1}M_{i}}\to\widehat{\pi_{1}M^{+}}$ is injective according to [41, Lemma 3.2.6]. Thus, we can view $\widehat{\pi_{1}M_{i}}$ and $\widehat{\pi_{1}N_{i}}$ as closed subgroups of $\widehat{\pi_{1}M^{+}}$ and $\widehat{\pi_{1}N^{+}}$ respectively. [51, Theorem 6.22] actually shows that $f^{+}$ sends each $\widehat{\pi_{1}M_{i}}$ to a conjugate of $\widehat{\pi_{1}N_{i}}$ in $\widehat{\pi_{1}N^{+}}$, and we denote $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ as $f^{+}$ composing with a conjugation and restricting on $\widehat{\pi_{1}M_{i}}$.
When $f$ respects the peripheral structure, we first claim that $f^{+}$ respects the peripheral structure. Indeed, when a boundary component $\partial_{k}M$ is compressible, the image of the homomorphism $\iota_{k_{\ast}}:\pi_{1}(\partial_{k}M)\to\pi_{1}(M)$ is isomorphic to $\mathbb{Z}$. Then $\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))}$ is the image of a continuous homomorphism ^$\widehat{\mathbb{Z}}\to\widehat{\pi_{1}M}$, and is hence procyclic. On the other hand, when $\partial_{k}M$ is incompressible, Proposition 2.3 implies that $\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))}\cong\widehat{\mathbb{Z}}^{2}$, which is not procyclic. The similar conclusion also holds for $N$. Since $f$ respects the peripheral structure, there is a one-to-one correspondence between the boundary components of $M$ and $N$, denoted by $\partial_{k}M\leftrightarrow\partial_{\sigma(k)}N$ such that $f(\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))})=g_{k}\cdot\overline{\rho_{\sigma(k)_{\ast}}(\pi_{1}(\partial_{\sigma(k)}M))}\cdot g_{k}^{-1}$ for some $g_{k}\in\widehat{\pi_{1}N}$. In particular, $\overline{\iota_{k_{\ast}}(\pi_{1}(\partial_{k}M))}\cong\overline{\rho_{\sigma(k)_{\ast}}(\pi_{1}(\partial_{\sigma(k)}M))}$. Distinguished by whether this group is procyclic or not, the above reasoning implies that $\partial_{\sigma(k)}N$ is incompressible if and only if $\partial_{k}M$ is incompressible. Note that the boundary of $M^{+}$ and $N^{+}$ consist of the incompressible boundary components of $M$ and $N$ respectively. For any incompressible boundary component $\partial_{k}M$ of $M$, $\iota_{k_{\ast}}^{+}:\pi_{1}(\partial_{k}M)\xrightarrow{\iota_{k_{\ast}}}\pi_{1}M\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}M^{+}$ is exactly the homomorphism induced by inclusion; and similarly, $\rho_{\sigma(k)_{\ast}}^{+}:\pi_{1}(\partial_{\sigma(k)}N)\xrightarrow{\rho_{\sigma(k)_{\ast}}}\pi_{1}N\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\pi_{1}N^{+}$ is induced by inclusion. Therefore, $f^{+}(\overline{\iota_{k_{\ast}}^{+}(\pi_{1}(\partial_{k}M)})=g_{k}^{+}\cdot\overline{\rho^{+}_{\sigma(k)_{\ast}}(\pi_{1}(\partial_{\sigma(k)}M))}\cdot(g_{k}^{+})^{-1}$, where $g_{k}^{+}$ denotes the image of $g_{k}$ in $\widehat{\pi_{1}N^{+}}$. This proves that $f^{+}$ respects the peripheral structure, with $\partial_{k}M\leftrightarrow\partial_{\sigma(k)}N$ for $\partial_{k}M$ incompressible.
Finally, it suffices to show that when $f^{+}$ respects the peripheral structure, each $f_{i}$ also respects the peripheral structure. Let $\partial_{k}M_{i}$ be a boundary component of $M_{i}$. Then the assumption that $f^{+}$ respects the peripheral structure implies that $f^{+}(\overline{\pi_{1}\partial_{k}M_{i}})$ is conjugate to a peripheral subgroup $\overline{\pi_{1}\partial_{l}N_{j}}\subseteq\widehat{\pi_{1}N_{j}}$ in $\widehat{\pi_{1}N^{+}}$. On the other hand, $f^{+}(\widehat{\pi_{1}M_{i}})$ is a conjugate of $\widehat{\pi_{1}N_{i}}$. Note that $\widehat{\pi_{1}N_{1}},\cdots,\widehat{\pi_{1}N_{s^{\prime}}}$ are factors of a free profinite product $\widehat{\pi_{1}N}=\widehat{\pi_{1}N_{1}}\amalg\cdots\amalg\widehat{\pi_{1}N_{s^{\prime}}}\amalg\widehat{F_{r_{1}^{\prime}+r_{2}^{\prime}}}$, so any conjugate of $\widehat{\pi_{1}N_{i}}$ intersects trivially $\widehat{\pi_{1}N_{j}}$ unless $i=j$. Therefore, $i=j$, and $f_{i}$ sends $\overline{\pi_{1}\partial_{k}M_{i}}$ to a conjugate of the closure of a peripheral subgroup in $\widehat{\pi_{1}N_{i}}$. Similarly, $f_{i}^{-1}$ sends each peripheral subgroup of $\widehat{\pi_{1}N_{i}}$ to a conjugate of a peripheral subgroup in $\widehat{\pi_{1}M_{i}}$. Indeed, this matches up a one-to-one correspondence between the boundary components of $M_{i}$ and $N_{i}$ according to [57, Proposition 6.7], which proves that $f_{i}$ respects the peripheral structure. ∎
@@
When $M$ has finite fundamental group, $\widehat{\pi_{1}N}\cong\widehat{\pi_{1}M}$ is finite, and the residual finiteness of $\pi_{1}N$ (Proposition 2.1) implies that $\pi_{1}N$ is also finite, so $N$ is also spherical. ∎
@@ 5.2. Detecting toroidality
We denote as in (5.2) the prime decomposition of $M$ and $N$. Then Lemma 5.1 implies that $s=s^{\prime}$, and $f$ induces isomorphisms $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ respecting the peripheral structure.
@@
According to the hyperbolization theorem [3, Theorem 1.13], $M_{i}$ is toroidal if and only if the following three conditions hold simultaneously.
@@
    $M_{i}$ is not a closed small Seifert fibered space, i.e. $M_{i}$ does not admit a Seifert fiberation over a sphere with at most 3 singular points;
@@
Moving on to condition (ii), when $M_{i}$ is a closed small Seifert fibered space, $N_{i}$ is also closed, and it follows from [54, Theorem 8.4] that $N_{i}$ is also a Seifert fibered space. When $\pi_{1}M_{i}$ is finite, the residual finiteness of $\pi_{1}N_{i}$ (cf. Proposition 2.1) implies that $\pi_{1}N_{i}$ is also finite, and $N_{i}$ is hence a small Seifert fibered space admitting $\mathbb{S}^{3}$ geometry. When $\pi_{1}M_{i}$ is infinite, [47, Theorem 6.1] implies that either $M_{i}$ and $N_{i}$ are homeomorphic, or $M_{i}$ and $N_{i}$ both admit $\mathbb{H}^{2}\times\mathbb{R}$ geometry and share the same base orbifold. In either case, $N_{i}$ is also a closed small Seifert fibered space.
For (iii), when $M_{i}$ is a Seifert fibered space with incompressible boundary, $M_{i}$ is irreducible and Lemma 5.1 implies that $N_{i}$ is also irreducible and boundary-incompressible. It then follows from [57, Theorem 5.6] that $N_{i}$ is also Seifert fibered. Therefore, $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ is a profinite isomorphism between Seifert fibered spaces with boundary, which respects the peripheral structure. [47, Theorem 6.7] then implies that $M_{i}$ and $N_{i}$ have the same base orbifolds. The list of Seifert fibered spaces in (iii) are only characterized by their base orbifolds. Therefore, if $M_{i}$ belongs to this list, then so does $N_{i}$, which finishes the proof. ∎
@@ 5.3. Detection of embedded Klein bottles
    $M$ is not a closed $Sol$-manifold, and the JSJ-decomposition [^1] We allow trivial JSJ-decomposition. of $M$ contains a Seifert fibered piece whose (one possible) base orbifold is
@@
The “if” direction is clear. Indeed, if $M$ satisfies condition (i), then $M$ is a torus semi-bundle, so $M$ contains an embedded submanifold homeomorphic to $\mathcal{N}$, and thus contains an embedded Klein bottle. If $M$ satisfies condition (ii), then $M$ contains a Seifert fibered submanifold, whose base orbifold is either a Möbius band, or a disk with two singular points of index $2$. Then this Seifert fibered submanifold is homeomorphic to $\mathcal{N}$, and thus $M$ contains an embedded Klein bottle.
Now we prove the “only if” direction. When $M$ has compressible boundary, since $M$ is aspherical, $M\cong D^{2}\times S^{1}$ and does not contain an embedded Klein bottle. Thus, we may assume that $M$ is boundary-incompressible. Suppose that $M$ contains an embedded Klein bottle $K^{2}$. Then the regular neighbourhood of $K^{2}$ in $M$ is homeomorphic to $\mathcal{N}$.
Let $T^{2}_{\mathcal{N}}$ denote the boundary torus of $\mathcal{N}$. We first claim that $T^{2}_{\mathcal{N}}$ is incompressible in $M$. Suppose $T^{2}_{\mathcal{N}}$ is compressible; since $M$ is aspherical, $T^{2}_{\mathcal{N}}$ bounds a solid torus in $M$ according to the loop theorem. Then $M=\mathcal{N}\cup_{T^{2}_{\mathcal{N}}}D^{2}\times S^{1}$. Note that $\mathcal{N}$ admits a two-fold cover $\widetilde{\mathcal{N}}\cong S^{1}\times S^{1}\times I$, where the covering map restricts to a homeomorphism from each boundary component of $\widetilde{\mathcal{N}}$ to the boundary of $\mathcal{N}$. Therefore, $M$ admits a two-fold cover $\widetilde{M}\cong D^{2}\times S^{1}\cup_{T^{2}}D^{2}\times S^{1}$, which is either $S^{2}\times S^{1}$ or a lens space. In each case, $\widetilde{M}$ is non-aspherical, so $M$ is non-aspherical, which contradicts with our assumption.
@@
Let us first consider two special cases. If $M^{\prime}\cong S^{1}\times S^{1}\times I$, then $M\cong\mathcal{N}$ and satisfies condition (ii). If $M^{\prime}\cong\mathcal{N}$, then $M$ is a torus semi-bundle. Thus, $M$ is either a $Sol$-manifold satisfying condition (i), or a Seifert fibered space whose base orbifold belongs to the following three possibilities:
  • $S^{2}$ with four singular points of index $2$;
  • $\mathbb{RP}^{2}$ with two singular points of index $2$;
  • a Klein bottle $K^{2}$.
Any one of these possibilities satisfies condition (ii).
Now we assume that $M^{\prime}$ is neither homeomorphic to $S^{1}\times S^{1}\times I$, nor to $\mathcal{N}$. Consider the JSJ-decomposition of $M^{\prime}$, whose collection of JSJ-tori is denoted by $\mathcal{T}^{\prime}$. Then there is a unique JSJ-piece in $M^{\prime}$ which contains the boundary component $T^{2}_{\mathcal{N}}$, and we denote this JSJ-piece as $M_{0}$.
When $M_{0}$ is a hyperbolic piece, let $\mathcal{T}=\mathcal{T}^{\prime}\cup\{T^{2}_{\mathcal{N}}\}$. Then [3, Proposition 1.9] implies that $\mathcal{T}$ is exactly the collection of JSJ-tori in $M$. Therefore, the JSJ-decomposition of $M$ contains $\mathcal{N}$ as a Seifert fibered piece, so it satisfies condition (ii).
When $M_{0}$ is a Seifert fibered piece, note that $M^{\prime}$ is not $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, so the base orbifold of $M_{0}$ has negative Euler characteristic, and $M_{0}$ admits a unique Seifert fiberation. On the other hand, $\mathcal{N}$ admits two Seifert fiberations, one of which has base orbifold a Möbius band, and the other has base orbifold a disk with two singular points of index $2$.
If the regular fiber of $M_{0}$ is not parallel to any of the two regular fibers of $\mathcal{N}$ on the common boundary $T^{2}_{\mathcal{N}}$, then by [3, Proposition 1.9], $\mathcal{T}=\mathcal{T}^{\prime}\cup\{T^{2}_{\mathcal{N}}\}$ is still the JSJ-tori of $M$. Similarly, $M$ contains $\mathcal{N}$ as a JSJ-piece, and satisfies condition (ii).
@@
Let $M$ and $N$ be orientable, aspherical Seifert fibered spaces with either empty or incompressible boundary, such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Suppose $M$ admits a Seifert fiberation over a base orbifold $\mathcal{O}$, where $\mathcal{O}$ is either non-orientable, or contains two singular points with index $2$. Then $N$ also admits a Seifert fiberation over a base orbifold $\mathcal{O}^{\prime}$, which is either non-orientable, or contains two singular points with index $2$.
@@
We first consider the case when $M$ is closed. Then [40, Corollary 4.12] implies that $N$ is also closed. According to [47, Theorem 6.1], if $M$ admits $\mathbb{E}^{3}$, $Nil$ or $\widetilde{\mathrm{SL}(2,\mathbb{R})}$ geometry, then $N$ is homeomorphic to $M$, and the conclusion is trivial. On the other hand, if $M$ admits $\mathbb{H}^{2}\times\mathbb{R}$ geometry, then $M$ admits a unique Seifert fiberation. [47, Theorem 6.2] implies that $N$ admits the same geometry, and has the same base orbifold as $M$, even though $M$ and $N$ may not be homeomorphic, which implies our conclusion.
Now let us consider the case when $M$ has non-empty boundary. Then the above reasoning shows that $N$ also has non-empty boundary. If $M$ is homeomorphic to $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, then it is easy to verify that $N\cong M$ (cf. [57, Proposition 8.2]), and the conclusion follows directly. Thus, we may assume that $M$ fibers over a hyperbolic orbifold. In this case, $M$ admits a unique Seifert fiberation, and we denote the base orbifold as $\mathcal{O}_{M}$. In addition, $N$ also fibers over a hyperbolic orbifold, and we denote the unique base orbifold as $\mathcal{O}_{N}$. In this case, the base orbifolds $\mathcal{O}_{M}$ and $\mathcal{O}_{N}$ may not be identical. However, in [57, Proposition 8.2], it is proven that the isomorphism type of $\widehat{\pi_{1}M}$ (resp. $\widehat{\pi_{1}N}$) determines whether the base orbifold $\mathcal{O}_{M}$ (resp. $\mathcal{O}_{N}$) is orientable, and determines the isomorphism type of the orbifold fundamental group $\pi_{1}(\mathcal{O}_{M})$ (resp. $\pi_{1}(\mathcal{O}_{N})$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies that $\mathcal{O}_{M}$ is non-orientable if and only if $\mathcal{O}_{N}$ is non-orientable. In addition, there is a one-to-one correspondence between the conjugacy classes of maximal torsion subgroups in the orbifold fundamental group with the singular points of a 2-orbifold [^2] We only allow discrete singular set in a 2-orbifold. with boundary. Therefore, $\mathcal{O}_{M}$ contains two singular points of index $2$ if and only if $\pi_{1}(\mathcal{O}_{M})\cong\pi_{1}(\mathcal{O}_{N})$ contains two conjugacy classes of maximal torsion subgroups isomorphic to $\mathbb{Z}/2\mathbb{Z}$, which implies that $\mathcal{O}_{N}$ also contains two singular points of index $2$. ∎
@@
Since $M$ contains an embedded Klein bottle, $M$ is not homeomorphic to $D^{2}\times S^{1}$ or $S^{2}\times S^{1}$. It follows from Corollary 5.3 that $N$ is also aspherical. In particular, $M$ is boundary-incompressible, and it follows from Lemma 5.1 that $N$ is also boundary-incompressible.

### v3

@@ 5.1. Detecting prime decomposition and asphericity
Let $M$ and $N$ be compact, orientable 3-manifolds with empty or toral boundary. The prime decomposition of $M$ and $N$ yields the following connected sums
@@
We Dehn fill the compressible boundary components of $M$ along their null-homotopic slopes, and obtain a boundary-incompressible 3-manifold
@@
To be explicit, up to a choice of conjugacy representatives, we have $\pi_{1}M^{+}=\pi_{1}M_{1}\ast\cdots\ast\pi_{1}M_{s}\ast F_{r_{1}+r_{2}}$ and $\widehat{\pi_{1}M^{+}}=\widehat{\pi_{1}M_{1}}\amalg\cdots\amalg\widehat{\pi_{1}M_{s}}\amalg\widehat{F_{r_{1}+r_{2}}}$, and likewise for $\widehat{\pi_{1}N^{+}}$. This can be checked by examining the universal property of free profinite products. With this decomposition, [51, Theorem 6.22] actually shows that $f^{+}$ sends each $\widehat{\pi_{1}M_{i}}$ to a conjugate of $\widehat{\pi_{1}N_{i}}$ in $\widehat{\pi_{1}N^{+}}$, and we denote $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ as $f^{+}$ composing with a conjugation and then restricted on $\widehat{\pi_{1}M_{i}}$.
When $f$ respects the peripheral structure, we first claim that $f^{+}$ respects the peripheral structure. Indeed, the closure of a peripheral subgroup corresponding to a compressible boundary component is isomorphic to $\widehat{\mathbb{Z}}$, while the closure of a peripheral subgroup corresponding to an incompressible boundary component is isomorphic to $\widehat{\mathbb{Z}}^{2}$. By Lemma 2.1, $\widehat{\mathbb{Z}}$ and $\widehat{\mathbb{Z}}^{2}$ are not isomorphic, so the bijection between the boundary components of $M$ and $N$ carried by $f$ distinguishes between the compressible ones and the incompressible ones. Note that the peripheral subgroups of $M^{+}$ and $N^{+}$ are exactly the images in $\pi_{1}M^{+}$ and $\pi_{1}N^{+}$ of the peripheral subgroups corresponding to the incompressible boundary components in $\pi_{1}M$ and $\pi_{1}N$. Then, by the same reasoning as the last paragraph in the proof of Theorem A, we derive that $f^{+}$ respects the peripheral structures.
[…]
@@
When $M$ has finite fundamental group, $\widehat{\pi_{1}N}\cong\widehat{\pi_{1}M}$ is finite, and the residual finiteness of $\pi_{1}N$ (Proposition 2.2) implies that $\pi_{1}N$ is also finite, so $N$ is also non-aspherical. ∎
@@ 5.2. Detecting toroidality
In this paper, the term toroidal is taken to mean geometrically toroidal.
Definition 5.4.
A compact, orientable 3-manifold $M$ is toroidal if $M$ contains an embedded incompressible torus which is not parallel to any boundary component.
@@
We denote as in (5.2) the prime decomposition of $M$ and $N$. Then, Lemma 5.1 implies that $s=s^{\prime}$, and $f$ induces isomorphisms $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ respecting the peripheral structure.
@@
Based on the hyperbolization theorem [3, Theorem 1.7.5], $M_{i}$ is toroidal if and only if the following three conditions hold simultaneously.
@@
    $M_{i}$ is not a closed small Seifert fibered space, i.e. $M_{i}$ does not admit a Seifert fibration over a sphere with at most 3 singular points;
@@
Moving on to condition (ii), it suffices to show that when $M_{i}$ is a closed small Seifert fibered space, $N_{i}$ is also a closed small Seifert fibered space. Indeed, $N_{i}$ is also closed since $f_{i}$ respects the peripheral structure, and it follows from [53, Theorem 8.4] that $N_{i}$ is also a Seifert fibered space. When $\pi_{1}M_{i}$ is finite, the residual finiteness of $\pi_{1}N_{i}$ (cf. Proposition 2.2) implies that $\pi_{1}N_{i}$ is also finite, and $N_{i}$ is hence a small Seifert fibered space admitting $\mathbb{S}^{3}$ geometry. When $\pi_{1}M_{i}$ is infinite, [47, Theorem 1.2] implies that either $M_{i}$ and $N_{i}$ are homeomorphic, or $M_{i}$ and $N_{i}$ both admit $\mathbb{H}^{2}\times\mathbb{R}$ geometry and share the same base orbifold. In either case, $N_{i}$ is also a closed small Seifert fibered space.
For (iii), it suffices to show that when $M_{i}$ belongs to this list, $N_{i}$ also belongs to the list. Since $M_{i}$ and $N_{i}$ are irreducible with non-empty incompressible toral boundary, it follows from a combination of [55] and [54] (see also [56, Theorem 5.5]) that $N_{i}$ is also Seifert fibered. Therefore, $f_{i}:\widehat{\pi_{1}M_{i}}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\widehat{\pi_{1}N_{i}}$ is a profinite isomorphism between Seifert fibered spaces with boundary, which respects the peripheral structure. [47, Theorem 1.3] then implies that $M_{i}$ and $N_{i}$ have the same base orbifolds. The list of Seifert fibered spaces in (iii) are only characterized by their base orbifolds. Therefore, when $M_{i}$ belongs to this list, so does $N_{i}$, which finishes the proof. ∎
@@ 5.3. Detection of embedded Klein bottles
    $M$ is not a closed $Sol$-manifold, and the JSJ-decomposition (possibly trivial) of $M$ contains a Seifert fibered piece whose (one possible) base orbifold is
@@
The “if” direction is clear. Indeed, if $M$ satisfies condition (i), then $M$ is a torus semi-bundle. In this case, $M$ contains an embedded submanifold homeomorphic to $\mathcal{N}$, and thus contains an embedded Klein bottle. If $M$ satisfies condition (ii), then $M$ contains a Seifert fibered submanifold, whose base orbifold is either a Möbius band, or a disk with two singular points of index $2$. Then, this Seifert fibered submanifold is homeomorphic to $\mathcal{N}$, and thus $M$ contains an embedded Klein bottle.
Now we prove the “only if” direction. When $M$ has compressible boundary, we have $M\cong D^{2}\times S^{1}$ since $M$ is aspherical, and in this case, $M$ does not contain an embedded Klein bottle. Thus, we may assume that $M$ is boundary-incompressible. Suppose that $M$ contains an embedded Klein bottle $K^{2}$. Then the regular neighbourhood of $K^{2}$ in $M$ is homeomorphic to $\mathcal{N}$.
Let $T^{2}_{\mathcal{N}}$ denote the boundary torus of $\mathcal{N}$. We first claim that $T^{2}_{\mathcal{N}}$ is incompressible in $M$. Suppose by contrary that $T^{2}_{\mathcal{N}}$ is compressible; since $M$ is aspherical, $T^{2}_{\mathcal{N}}$ bounds a solid torus in $M$ according to the loop theorem. Then $M=\mathcal{N}\cup_{T^{2}_{\mathcal{N}}}D^{2}\times S^{1}$. Note that $\mathcal{N}$ admits a two-fold cover $\widetilde{\mathcal{N}}\cong S^{1}\times S^{1}\times I$, where the covering map restricts to a homeomorphism from each boundary component of $\widetilde{\mathcal{N}}$ to the boundary of $\mathcal{N}$. Therefore, $M$ admits a two-fold cover $\widetilde{M}\cong D^{2}\times S^{1}\cup_{T^{2}}D^{2}\times S^{1}$, which is either $S^{2}\times S^{1}$ or a lens space. In either case, $\widetilde{M}$ is non-aspherical, so $M$ is non-aspherical, which contradicts with our assumption.
@@
Let us first consider two special cases. If $M^{\prime}\cong S^{1}\times S^{1}\times I$, then $M\cong\mathcal{N}$ and satisfies condition (ii). If $M^{\prime}$ is homeomorphic to $\mathcal{N}$, then $M$ is a torus semi-bundle. Both $\mathcal{N}$ and $M^{\prime}$ admit two Seifert fibrations. Up to isotopy, if one of the two possible regular fibers of $\mathcal{N}$ is glued along $T^{2}_{\mathcal{N}}$ to one of the two possible regular fibers of $M^{\prime}$, then the Seifert fibration structures on both sides combine into a Seifert fibration of $M$ whose base orbifold is obtained by gluing the base orbifolds of both sides along their boundary circles. In particular, the base orbifold belongs to the following three possibilities:
  • $S^{2}$ with four singular points of index $2$,
  • $\mathbb{RP}^{2}$ with two singular points of index $2$,
  • a Klein bottle $K^{2}$,
any of which satisfies condition (ii). Otherwise, the gluing along $T^{2}_{\mathcal{N}}$ is never fiber-parallel, so by [3, Proposition 1.6.2], $T^{2}_{\mathcal{N}}$ is a JSJ-torus of $M$. In this case, $M$ is a $Sol$-manifold satisfying condition (i).
Now we assume that $M^{\prime}$ is neither homeomorphic to $S^{1}\times S^{1}\times I$, nor to $\mathcal{N}$, so $M^{\prime}$ is not a closed $Sol$-manifold. Consider the JSJ-decomposition of $M^{\prime}$, whose collection of JSJ-tori is denoted by $\mathcal{T}^{\prime}$ (possibly empty). Then, there is a unique JSJ-piece in $M^{\prime}$ which contains the boundary component $T^{2}_{\mathcal{N}}$, and we denote this JSJ-piece as $M_{0}$.
When $M_{0}$ is a hyperbolic piece, let $\mathcal{T}=\mathcal{T}^{\prime}\cup\{T^{2}_{\mathcal{N}}\}$. Then [3, Proposition 1.6.2] implies that $\mathcal{T}$ is exactly the collection of JSJ-tori in $M$. Therefore, the JSJ-decomposition of $M$ contains $\mathcal{N}$ as a Seifert fibered piece, so it satisfies condition (ii).
When $M_{0}$ is a Seifert fibered piece, note that $M^{\prime}$ is not $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, so the base orbifold of $M_{0}$ has negative Euler characteristic, and $M_{0}$ admits a unique Seifert fibration. On the other hand, $\mathcal{N}$ admits two Seifert fibrations, one of which has base orbifold a Möbius band, and the other has base orbifold a disk with two singular points of index $2$.
If the regular fiber of $M_{0}$ is not parallel to any of the two regular fibers of $\mathcal{N}$ on the common boundary $T^{2}_{\mathcal{N}}$, then by [3, Proposition 1.6.2], $\mathcal{T}=\mathcal{T}^{\prime}\cup\{T^{2}_{\mathcal{N}}\}$ is still the JSJ-tori of $M$. Similarly, $M$ contains $\mathcal{N}$ as a JSJ-piece, and satisfies condition (ii).
@@
Let $M$ and $N$ be orientable, aspherical Seifert fibered spaces with either empty or incompressible boundary, such that $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$. Suppose $M$ admits a Seifert fibration over a base orbifold $\mathcal{O}$, where $\mathcal{O}$ is either non-orientable, or contains two singular points with index $2$. Then $N$ also admits a Seifert fibration over a base orbifold $\mathcal{O}^{\prime}$, which is either non-orientable, or contains two singular points with index $2$.
@@
We first consider the case when $M$ is closed. Then [40, Corollary 4.12] implies that $N$ is also closed. According to [47, Theorem 1.2], if $M$ admits $\mathbb{E}^{3}$, $Nil$ or $\widetilde{\mathrm{SL}(2,\mathbb{R})}$ geometry, then $N$ is homeomorphic to $M$, and the conclusion is trivial. On the other hand, if $M$ admits $\mathbb{H}^{2}\times\mathbb{R}$ geometry, then $M$ admits a unique Seifert fibration. [47, Theorem 5.2] implies that $N$ admits the same geometry, and has the same base orbifold as $M$ even though $M$ and $N$ may not be homeomorphic, which implies our conclusion.
Now let us consider the case when $M$ has non-empty boundary. Then the above reasoning shows that $N$ also has non-empty boundary. If $M$ is homeomorphic to $S^{1}\times S^{1}\times I$ or $\mathcal{N}$, then it is easy to verify that $N\cong M$ (cf. [56, Proposition 8.2]), and the conclusion follows directly. Thus, we may assume that $M$ fibers over a hyperbolic orbifold. In this case, $M$ admits a unique Seifert fibration, and we denote the base orbifold as $\mathcal{O}_{M}$. In addition, $N$ also fibers over a hyperbolic orbifold, and we denote the unique base orbifold as $\mathcal{O}_{N}$. In this case, the base orbifolds $\mathcal{O}_{M}$ and $\mathcal{O}_{N}$ may not be identical. However, in [56, Proposition 8.2], it is proven that the isomorphism type of $\widehat{\pi_{1}M}$ (resp. $\widehat{\pi_{1}N}$) determines whether the base orbifold $\mathcal{O}_{M}$ (resp. $\mathcal{O}_{N}$) is orientable, and determines the isomorphism type of the orbifold fundamental group $\pi_{1}(\mathcal{O}_{M})$ (resp. $\pi_{1}(\mathcal{O}_{N})$). Thus, $\widehat{\pi_{1}M}\cong\widehat{\pi_{1}N}$ implies that $\mathcal{O}_{M}$ is non-orientable if and only if $\mathcal{O}_{N}$ is non-orientable. In addition, there is a one-to-one correspondence between the conjugacy classes of maximal torsion subgroups in the orbifold fundamental group with the cone points of a compact 2-orbifold with discrete singular set and non-empty boundary. Therefore, $\mathcal{O}_{M}$ contains two singular points of index $2$ if and only if $\pi_{1}(\mathcal{O}_{M})\cong\pi_{1}(\mathcal{O}_{N})$ contains two conjugacy classes of maximal torsion subgroups isomorphic to $\mathbb{Z}/2\mathbb{Z}$, which implies that $\mathcal{O}_{N}$ also contains two singular points of index $2$. ∎
@@
Since $M$ contains an embedded Klein bottle, $M$ is not homeomorphic to $D^{2}\times S^{1}$. As $M$ is aspherical, it follows from Corollary 5.3 that $N$ is also aspherical. In addition, $M$ is boundary-incompressible, and it follows from Lemma 5.1 that $N$ is also boundary-incompressible.